OpenStudy (anonymous):

Let f(x) = 2 + 1/x. Which of the following are zeros of f(x)? *check all that apply* a. x = 0 b. x = -1/2 c. x = 1/2 d. none of the above

OpenStudy (agent0smith):

You need to find x, by setting f(x) equal to zero, like so \[\Large 2 + \frac{ 1 }{ x } = 0\] Try solving that for x.

OpenStudy (anonymous):

is it -2?

OpenStudy (agent0smith):

No... but start by subtracting 2 from both sides.

OpenStudy (agent0smith):

If you use x=-2, 2+(1/-2) = 3/2, which is not zero.

OpenStudy (anonymous):

So when I subtract 2 from both sides I get 1/x = -2

OpenStudy (agent0smith):

Correct. Now, you can cross-multiply and then find x. Write the -2 like so: \[\Large \frac{ 1 }{ x} = \frac{ -2 }{ 1}\] or you can just invert both fractions to find x.

OpenStudy (anonymous):

When I cross multiply should I get 1 = -2x?

OpenStudy (anonymous):

Or x = -2?

OpenStudy (agent0smith):

1 = -2x is correct. x=-2 is not.

OpenStudy (anonymous):

Thought so. So what do I do next?

OpenStudy (anonymous):

Divide?

OpenStudy (agent0smith):

Yes.

OpenStudy (anonymous):

1/-2? Or -1/2?

OpenStudy (anonymous):

So is it just b?

OpenStudy (agent0smith):

1/-2? Or -1/2 those are the same thing \[\Large \frac{ -1 }{ 2 } = \frac{ 1 }{ -2 } = -\frac{ 1 }{ 2 }\]